OpenStudy (anonymous):

change the equation into slope-intercept form of this problem: –3x + 2y = 8 Please don't just give me the answer, if somebody could show me the steps of how to answer this that would be awesome! Thank you :))

OpenStudy (mertsj):

Solve it for y

OpenStudy (mertsj):

First add 3x to both sides and then divide both sides by 2

OpenStudy (anonymous):

Slope intercept form is y=mx+b, So, –3x + 2y = 8 +3x +3x 2y=3x+8 /2 /2 \[y=\frac{ 3 }{ 2 }x+4\]

OpenStudy (anonymous):

Do you see how I did that? :o)

OpenStudy (anonymous):

I see what you did thar. Thank you so much! :)))

OpenStudy (anonymous):

@Butterfly16